# Extreme Compression of Large Language Models via Additive Quantization

### v2

@@ 4.3 Inference Speed
Although our primary objective is to maximize accuracy for a given model size, AQLM can also be practical in terms of inference latency. To demonstrate this, we implemented efficient GPU and CPU kernels for a few hardware-friendly configurations of AQLM. The results can be found in Table 4. For GPU inference, we targeted quantized Llama 2 models with 16-bit codebooks, corresponding to 2.07 bits for Llama 2 70B, 2.18 bits for 13B, and 2.29 bits for 7B models (see Table 1), as well as a 2x8-bit codebook model with perplexity 7.98 on Wiki2. For each model we benchmark the matrix-vector multiplication subroutine performance on a standard layer. The results show that AQLM can execute at speeds comparable to or better than FP16. End-to-end generative numbers on a preliminary HuggingFace integration can be found in Appendix H: for instance, we can achieve $\sim$ 6 tokens/s on Llama 2 70B in this setting. We observe that multiple smaller codebooks allow efficient GPU cache utilization, leading to greater speedup, at the price of slightly lower accuracy.
@@ 5 Conclusion and Future Work
We presented AQLM, a new form of additive quantization (AQ) targeted to LLM compression, which significantly improved the state-of-the-art results for LLM quantization in the regime of 2 and 3 bits per weight. In terms of limitations, AQLM is more computationally-expensive than direct post-training quantization methods, such as RTN or GPTQ, specifically because of the use of a more complex coding representation. Yet, despite the more sophisticated encoding and decoding, we have shown AQLM lends itself to efficient implementation on both CPU and GPU. Overall, we find it remarkable that, using AQLM, massive LLMs can be executed accurately and efficiently using little memory.
6 Acknowledgements
Authors would like to thank Ruslan Svirschevski for his help in solving technical issues with AQLM and baselines. We also thank Tim Dettmers for helpful discussions on the structure of weights in modern LLMs and size-accuracy trade-offs.The authors would also like to thank Daniil Pavlov for his assistance with CPU benchmarking. Finally, authors would like to thank the communities of ML enthusiasts known as LocalLLaMA [^3] https://www.reddit.com/r/LocalLLaMA/ and Petals community on discord [^4] https://github.com/bigscience-workshop/petals/ for the crowd wisdom about running LLMs on consumer devices.
Appendix A Full model fine-tuning
The block-wise finetuning procedure, introduced in 3.4, considerably improves performance of compressed models. However, block-wise finetuning optimizes the loss only at the level of a current transformer block and is agnostic of the actual task of interest. To minimize the target loss, one can run backpropagation through the whole model and optimize all trainable parameters to directly optimize the end objective.
@@ Appendix A Full model fine-tuning
One can minimize the error between the quantized model and the floating-point model on some calibration set. The parameters being optimized (namely the codebooks, scales and the non-quantized parameters) typically constitute a small fraction of the total number of parameters in the original model. Therefore, the proposed distillation method is an instance of parameter-efficient finetuning.
Standard Knowledge distillation (Hinton et al., 2015) optimizes the KL-divergence between the outputs of teacher and student models. However, according to the literature, (Sun et al., 2019; Kurtic et al., 2023) transfer of intermediate representation facilitates faster convergence and enhances performance. Hence, we decided to perform knowledge transfer between some intermediate activations in addition to the model output. Specifically, we take a subset of transformer blocks $\{l_{0},l_{1}\ldots,l_{K-1}\},l_{i}\in[0,L-1]$, where $L$ is the number of transformers block in the model, and optimize the following objective:
$\sum_{i=0}^{K-1}\mathcal{L}(x_{l_{i}}^{(s)},x_{l_{i}}^{(t)})$ (9)
Where $x_{l_{i}}^{(s)}$ and $x_{l_{i}}^{(t)}$ are the student and teacher output activations respectively at block $l_{i}$ and $\mathcal{L}(*,*)$ is some loss function. We considered the following options:
  • $L_{2}$ loss: $\|x-y\|_{2}^{2}$
  • Normalized $L_{2}$ loss: $\|x-y\|_{2}^{2}/\|y\|_{2}^{2}$
  • Cosine distance: $\frac{1}{2}\left(1-\frac{(x,y)}{\|x\|\|y\|}\right)$
We tune all models for a single epoch on the same calibration set used for additive quantization and block finetuning. We observe that longer training leads to overfitting. Learning rate of $\eta=1e-5$ and batch size of $32$ with Adam optimizer are used in full model finetuning experiments. We find that simple $L_{2}$ loss outperforms the alternatives (Table 5), and that the number of intermediate activations used in the loss function barely affects the final performance (Table 6).
Table 5: Ablation on the choice of loss for Knowledge distillation.
Loss Wiki2normal-$\downarrow$ C4normal-$\downarrow$
$L_{2}$ $\mathbf{6.57}$ $\mathbf{8.48}$
Normalized $L_{2}$ 6.63 8.57
Cosine distance 6.65 8.59
Table 6: Ablation on the choice of blocks subset for Knowledge distillation. Empty set means no knowledge distillation applied.
Blocks Wiki2normal-$\downarrow$ C4normal-$\downarrow$
[…]
$\{31\}$ $\mathbf{6.52}$ 8.51
$\{15,31\}$ 6.57 $\mathbf{8.48}$
$\{7,15,23,31\}$ 6.55 8.49
$\{3,7,11,15,19,23,27,31\}$ 6.57 8.54
Overall, the improvement due to knowledge distillation is not very significant while increasing the complexity of the quantization pipeline, therefore, we decided to not include KD in our main setup.
@@ Appendix B Code reproducibility
We share the code for our method in the GitHub repository https://github.com/vahe1994/AQLM. The hyperparameters for our experimental setup are discussed in Appendix C.
@@ Appendix C Experimental Configurations
Calibration set. All methods were calibrated on a slice of RedPajama-v1 dataset (Computer, 2023) for both Llama and Mistral/Mixtral family models. We used the same context length as models were trained on, for Llama-2 4096 and for Mistral/Mixtral 8192.
For Llama-2 experiments, we used 8M tokens as a calibration set for SpQR, GPTQ, and AQLM. Quip, however, was calibrated on 4M tokens due to OOM errors when trying to use more samples. Taking into account the fact that after 2M tokens improvement of methods results is fairly small we chose to report these numbers as is. For Quip#, we used Llama-2 and Mistral’s quantized models provided by authors in their GitHub repository. To the best of our knowledge, they used 6k samples for calibration with a context length of 4096/8192.
@@
Quip# For Llama-2 and Mistral models we used the officially published quantized models. For Mixtral we adapted the code to work with the model’s architecture and quantized it with the recommended set of parameters. For both AQLM and QuIP#, we don’t quantize gate linear layer in Mixtral, because it contains relatively small amount of paramters and have severe impact on performance.
@@
Appendix D Ablation analysis
@@
Name Wiki2normal-$\downarrow$ C4normal-$\downarrow$
@@
Number of samples. We verify our choice of calibration hyperparameters. Traditionally, most PTQ algorithms use several hundred calibration sequences (e.g. Frantar et al. (2022a) has 128). In our experiments, we evaluate both AQLM and baselines with additional calibration data. Our original motivation for that was to avoid potential overfitting when fine-tuning entire transformer blocks. To test this assumption, we run our algorithm with different calibration set sizes, varying from 128 to 4096 sequences. For each size, we report the average perplexity on WikiText2 over 3 runs, along with standard deviations. The results in Table 8 demonstrate that increasing the number of samples leads to gradual reduction in perplexity with seemingly diminishing returns. Since the perplexity is still monotonically improving from 128 to 4096 samples, it is possible that larger sample sizes would yield further improvements.
Number of codebooks vs groups. Finally, we conducted an additional set of experiments on LLama-2 7B models to see perplexity dependence on simultaneous change of both codebooks and groups keeping compression rate fixed to 2bits. 9
We can see that the simultaneous increase in both the number of codebooks and groups led to a decrease in average perplexity and stopped when the number of groups reached 64.
Table 8: Wikitext2 PPL as a function of calibration set size for Llama 2 (7B) quantized to 2.3 bits with AQLM, averaged over 3 runs. SD stands for adjusted standard deviation.
@@
Table 9: Wikitext2 PPL as a function of from groups and number of codebook for Llama 2 (7B) quantized with approximately 2 bits quantization.
# of codebooks Groups size Average PPL
1 4 41.79
2 8 11.96
4 16 8.53
8 32 7.83
15 64 7.86
Appendix E Additional experiments
In this section we report additional experimental results for Mixtral(Jiang et al., 2024) , Mistral7B(Jiang et al., 2023) and LLama-2 model.
Table 10: Evaluation of quantized Llama 2 models for 4+ bits per parameter. The table reports perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
Size Method Avg bits Wiki2normal-$\downarrow$ C4normal-$\downarrow$ WinoGrandenormal-$\uparrow$ PiQAnormal-$\uparrow$ HellaSwagnormal-$\uparrow$ ArcEnormal-$\uparrow$ ArcCnormal-$\uparrow$ Average accuracynormal-$\uparrow$
@@ E.1 Mixtral
Table 11: Evaluation of quantized Mixtral (Jiang et al., 2024) models for 3 and 4 bits per parameter. The table reports perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
Size Method Avg bits Wiki2normal-$\downarrow$ C4normal-$\downarrow$ WinoGrandenormal-$\uparrow$ PiQAnormal-$\uparrow$ HellaSwagnormal-$\uparrow$ ArcEnormal-$\uparrow$ ArcCnormal-$\uparrow$ Average accuracynormal-$\uparrow$
@@ E.2 Llama-2
Table 12: Wikitext2 and C4 PPL for quantized Llama 2 models for 2x8 codebooks models.
Name Wiki2normal-$\downarrow$ C4normal-$\downarrow$
7B 7.98 10.38
70B 4.83 6.66
Table 13: Evaluation of quantized Mistral7B (Jiang et al., 2023) models for 2, 3 and 4 bits per parameter: perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
Size Method Avg bits Wiki2normal-$\downarrow$ C4normal-$\downarrow$ WinoGrandenormal-$\uparrow$ PiQAnormal-$\uparrow$ HellaSwagnormal-$\uparrow$ ArcEnormal-$\uparrow$ ArcCnormal-$\uparrow$ Average accuracynormal-$\uparrow$
@@
E.3 Mistral
Finally we evaluate AQLM and QuIP# quantization on Mistral7b (Jiang et al., 2023) model for 3 and 4 bits in Table 13. In 2bits, QuIP# slightly outperform AQLM on most benchmarks. And for 4bits setup results are very close across the board.
Appendix F Pareto optimality
We visualize WikiText perplexity of Llama-2 7B,13B,70B models quantized with AQLM and QuIP# as plotted against quantized weight size in bytes and report it in Figure 5. Our method outperforms QuIP# in terms of perplexity in WikiText2 across all model sizes.
Additionally, in Figure 6, we show perplexity on WikiText for AQLM method against size of quantized parameters. We can notice that starting around 3.7 GiB of quantized weights, which correspond to 2.5 bits compression on Llama-2 13B model, it is more advantageous to compress 13B model rather 7B model at the same model size in bytes.
[Uncaptioned image] Figure 5: Comparison of AQLM relative to QuIP# on LLAMA-2 7B, 13B, and 70B models.
[Uncaptioned image] Figure 6: Model optimality for AQLM on LLAMA 2 7, 13, and 70B models.
Appendix G Estimating model size

### v3

@@ 4.4 Inference Speed
Although our primary objective is to maximize accuracy for a given model size, AQLM can also be practical in terms of inference latency. To demonstrate this, we implemented efficient GPU and CPU kernels for a few hardware-friendly configurations of AQLM. The results can be found in Table 5. For GPU inference, we targeted quantized Llama 2 models with 16-bit codebooks, corresponding to 2.07 bits for Llama 2 70B, 2.19 bits for 13B, and 2.29 bits for 7B models (see Table 1, 4), as well as a 2x8-bit codebook model with perplexity 6.57 on Wiki2(see Table 12). For each model we benchmark the matrix-vector multiplication subroutine performance on a standard layer. The results show that AQLM can execute at speeds comparable to or better than FP16. End-to-end generative numbers with HuggingFace integration can be found in Appendix I: for instance, we can achieve ${\approx}14$ tokens/s on Llama 2 70B in this setting. We observe that multiple smaller codebooks allow efficient GPU cache utilization, leading to greater speedup, at the price of slightly lower accuracy.
@@ 5 Conclusion and Future Work
We presented AQLM, a new form of additive quantization (AQ) targeting LLM compression, which significantly improved the state-of-the-art results for LLM quantization in the regime of 2 and 3 bits per weight. In terms of limitations, AQLM is more computationally-expensive than direct post-training quantization methods, such as RTN or GPTQ, specifically because of the use of a more complex coding representation. Yet, despite the more sophisticated encoding and decoding, we have shown AQLM lends itself to efficient implementation on both CPU and GPU. Overall, we find it remarkable that, using AQLM, massive LLMs can be executed accurately and efficiently using little memory.
While AQLM already achieves substantial improvements in low-bit quantization, there are several promising directions for further improvement that we did not explore in this work. One such direction is better fine-tuning strategies. In Section 4.2 we found that better fine-tuning algorithms (Tseng et al., 2024; Malinovskii et al., 2024) can significantly improve quantized model accuracy. We believe that AQLM can benefit from a more systematic exploration of fine-tuning algorithms in future work. Another promising direction is generalizing AQLM to other quantization scenarios. While our work is focused around LLM quantization, the underlying algorithm can potentially be adapted to other problems, e.g. quantizing computer vision models, compressing LLM attention caches for long sequences, and others.
Acknowledgements
Authors would like to thank Ruslan Svirschevski for his help in solving technical issues with AQLM and baselines. We also thank Tim Dettmers for helpful discussions on the structure of weights in modern LLMs and size-accuracy trade-offs. The authors would also like to thank Daniil Pavlov for his assistance with CPU benchmarking. Finally, authors would like to thank the communities of ML enthusiasts known as LocalLLaMA [^5] https://www.reddit.com/r/LocalLLaMA/ and Petals community on discord [^6] https://github.com/bigscience-workshop/petals/ for the crowd wisdom about running LLMs on consumer devices. Egiazarian Vage and Denis Kuznedelev and Andrei Panferov were supported by the grant for research centers in the field of AI provided by the Analytical Center for the Government of the Russian Federation (ACRF) in accordance with the agreement on the provision of subsidies (identifier of the agreement  [^00] D730321P5Q0002) and the agreement with HSE University No. 70-2021-00139
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A End-to-end fine-tuning
The block-wise finetuning procedure, introduced in 3.4, considerably improves performance of compressed models. However, block-wise finetuning optimizes the loss only at the level of a current transformer block and is agnostic of the actual task of interest. To minimize the target loss, one can run backpropagation through the whole model and directly optimize all trainable parameters to minimize a model-level objective function.
@@ Appendix A End-to-end fine-tuning
One can minimize the error between the quantized model and the floating-point model on some calibration set. The parameters being optimized (namely the codebooks, scales and the non-quantized parameters) typically constitute a small fraction of the total number of parameters in the original model. Therefore, the proposed distillation method resembles parameter-efficient finetuning (PEFT) in both optimization and memory footprint.
To transfer the knowledge from the original model to the quantized one, we adopt Knowledge Distillation (Hinton et al., 2015) where the student model is taught to mimic the output of a teacher given the same input. We follow the setup from QuIP# (Tseng et al., 2024) that uses KL divergence between the outputs of teacher and student models:
$\mathcal{L}=\frac{1}{N}\sum_{i=0}^{N-1}D_{KL}(p_{s}(\mathbf{x}_{i}),p_{t}(\mathbf{x}_{i}))$ (9)
Above $D_{KL}$ is the Kullback–Leibler divergence and $p_{s}$, $p_{t}$ are the student and teacher probabilities given input sequence $\mathbf{x}_{i}$.
Despite its simplicity, this fine-tuning procedure often significantly improves performance of the compressed model.
We fine-tune all models on $4{-}16$M training tokens: $1{-}4$k sequences of length 4k for Llama 2 models (Touvron et al., 2023) and 512 sequences of length 8k for Mixtral (Jiang et al., 2024). We fine-tune on the same data as during initial calibration (i.e. samples from RedPajama (Computer, 2023)) and use Adam (Kingma & Ba, 2015) optimizer with constant learning rate $10^{-5}$ without weight decay. Batch size is set to $8{-}16$ sequences. A single epoch of fine-tuning turns out to be sufficient, and longer training leads to marginal improvements.
@@ Appendix B Code reproducibility
We share the code for our method in the GitHub repository https://github.com/Vahe1994/AQLM/tree/AQLM_camera_ready. The hyperparameters for our experimental setup are discussed in Appendix C.
@@ Appendix C Experimental Configurations
Calibration set. All methods were calibrated on a slice of RedPajama-v1 dataset (Computer, 2023) for both Llama and Mistral/Mixtral family models. We used the same context length as models were trained on, for Llama 2 4096 and for Mistral/Mixtral 8192.
For Llama 2 experiments, we used 8M tokens as a calibration set for SpQR, GPTQ, and AQLM. Quip, however, was calibrated on 4M tokens due to OOM errors when trying to use more samples. Taking into account the fact that after 2M tokens improvement of methods results is fairly small we chose to report these numbers as is. For Quip#, we used Llama 2 and Mistral’s quantized models provided by authors in their GitHub repository. To the best of our knowledge, they used 6k samples for calibration with a context length of 4096/8192.
@@
Quip# For Llama 2 and Mistral models we used the officially published quantized models. For Mixtral we adapted the code to work with the model’s architecture and quantized it with the recommended set of parameters. For both AQLM and QuIP#, we don’t quantize gate linear layer in Mixtral, because it contains relatively small amount of paramters and have severe impact on performance.
@@
Hyperparameters for end-end fine-tuning discussed at the end of Appendix A.
Appendix D Quantization time
AQLM quantization takes considerably longer to calibrate than simpler quantization methods such as RTN or GPTQ. This only impacts quantization time, not inference time.
Quantizing a 7B model with default configuration takes about 1 day on a single A100 gpu. Similarly, quantizing a 70B model on a single GPU would take 10-14 days. However, the procedure can be parallelized across multiple GPU: 7B quantization takes  14h on 2 GPUs, and 70B quantization takes 3-4 days on 8 GPUs.
Full model fine-tuning with default configuration for 7B model would take 3-6 hours on four A100 , for 13B 10-16 hours on four A100, and for 70B 1-2 days on 8 A100.
Finally, the quantization time is dependent on the quantization configuration and its hyperparameters. Tweaking these parameters, e.g. by reducing the number of beams, can achieve notable speedups of 2-4x during quantization, but at the cost of lower model accuracy.
Table 6: Evaluation of quantized Llama 2 end-to-end fine-tuned models for 3-3.1 bits per parameter, with the same metrics as in Table 1.
[…]
7B 16 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 3.04 5.38 7.01 65.35 77.31 55.49 66.79 38.48 60.68
QuIP# 3.04 5.41 7.04 66.85 77.31 55.32 68.43 38.99 61.38
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 3.03 4.78 6.33 68.75 78.45 58.54 72.94 42.75 64.29
QuIP# 3.01 4.78 6.35 68.03 77.86 57.56 72.18 41.38 63.40
70B 16 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 3.01 3.36 5.17 75.30 80.69 63.48 77.99 50.26 69.54
QuIP# 3.00 3.35 5.15 76.40 81.45 63.53 77.53 50.77 69.94
Appendix E Ablation analysis
@@
Name Wiki2$\downarrow$ C4$\downarrow$
@@
Number of samples. We verify our choice of calibration hyperparameters. Traditionally, most PTQ algorithms use several hundred calibration sequences (e.g. Frantar et al. (2022a) has 128). In our experiments, we evaluate both AQLM and baselines with additional calibration data. Our original motivation for that was to avoid potential overfitting when fine-tuning entire transformer blocks. To test this assumption, we run our algorithm with different calibration set sizes, varying from 128 to 4096 sequences. For each size, we report the average perplexity on WikiText-2 over 3 runs, along with standard deviations. The results in Table 8 demonstrate that increasing the number of samples leads to gradual reduction in perplexity with seemingly diminishing returns. Since the perplexity is still monotonically improving from 128 to 4096 samples, it is possible that larger sample sizes would yield further improvements.
Number of codebooks vs groups. Finally, we conducted an additional set of experiments on LLama 2 7B models to see perplexity dependence on simultaneous change on WikiText-2 of both codebooks and groups keeping compression rate fixed to 2bits. We present both AQLM with and without end-to-end fine-tuning in Table 9.
Table 8: WikiText-2 PPL as a function of calibration set size for Llama 2 (7B) quantized to 2.3 bits with AQLM, averaged over 3 runs. SD stands for adjusted standard deviation.
@@
Table 9: WikiText-2 PPL as a function of from groups and number of codebook for Llama 2 (7B) quantized with approximately 2 bits quantization.
Method Setup Average PPL
AQLM 2x8gs8 7.6107
4x8gs16 8.1394
8x8gs32 7.3755
15x8gs64 7.8459
AQLM 2x8gs8 6.5746
8x8gs32 6.6126
15x8gs64 6.6602
Appendix F Additional experiments
In this section we report additional experimental results for Mixtral(Jiang et al., 2024), Mistral7B(Jiang et al., 2023) and LLama 2 model.
Table 10: Evaluation of quantized Llama 2 models for 4+ bits per parameter. The table reports perplexity on WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
[…]
@@ F.1 Mixtral
Table 11: Evaluation of quantized Mixtral (Jiang et al., 2024) models for 3 and 4 bits per parameter. The table reports perplexity on WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. The primary metrics are Wiki2 (PPL, lower is better), C4 (PPL, lower is better) and Average accuracy (percentage, higher is better).
[…]
@@ F.2 Llama 2
Table 12: Evaluation of quantized Llama 2 for 2x8groupsize8 codebooks models. We report perplexity on WikiText-2 (Merity et al., 2016) & C4 (Raffel et al., 2020) and accuracy for zero-shot tasks. The Average accuracy is the mean of 5 zero-shot tasks. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
[…]
7B 16 5.12 6.63 67.25 78.40 56.67 69.36 39.51 62.24
AQLM 2 7.61 9.68 62.27 71.87 46.41 61.03 30.03 54.32
AQLM 2 6.57 8.60 63.22 74.54 50.08 61.28 31.83 56.19
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 2 6.54 8.77 55.96 71.06 48.29 62.50 31.40 53.84
AQLM 2 5.63 7.55 6385 77.04 54.19 67.85 37.20 60.03
70B 16 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 2 4.21 5.99 73.48 79.54 61.29 74.49 46.84 67.13
Table 13: Evaluation of quantized Mistral7B (Jiang et al., 2023) models for 2, 3 and 4 bits per parameter: perplexity on WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
[…]
@@
F.3 Mistral
Finally, we evaluate AQLM and QuIP# quantization on Mistral 7b (Jiang et al., 2023) model for 3 and 4 bits in Table 13. In 2 bits, QuIP# slightly outperform AQLM on most benchmarks. And for 4 bits setup results are very close across the board.
Appendix G Pareto optimality
We visualize WikiText-2 perplexity of Llama-2 7B, 13B, 70B models quantized with AQLM and QuIP# as plotted against quantized weight size in bytes and report it in Figure 5. Our method outperforms QuIP# in terms of perplexity in WikiText-2 across all model sizes.
Additionally, in Figure 6, we show perplexity on WikiText-2 for AQLM method against size of quantized parameters. We can notice that starting around 3.7 GiB of quantized weights, which correspond to 2.5 bits compression on Llama 2 13B model, it is more advantageous to compress 13B model rather 7B model at the same model size in bytes.
Figure 5: Comparison of AQLM relative to QuIP# on Llama 2 7B, 13B, and 70B models.
Figure 6: Model optimality for AQLM on Llama 2 7, 13, and 70B models.
Appendix H Estimating model size